OpenStudy (anonymous):

what is the standard form equation of the line y=x/5-2?

OpenStudy (unklerhaukus):

do you mean \[y=mx+b\] or \[ax+by+c=0\]

OpenStudy (anonymous):

i think it ax+by+c=0

OpenStudy (anonymous):

y= x/3 I be thinkin

OpenStudy (anonymous):

that is not the correct answer.

OpenStudy (anonymous):

Do you have something better ;D Ax=By=C anyway http://www.learningwave.com/lwonline/algebra_section2/equation1.html

OpenStudy (anonymous):

*Ax+By=C

Directrix (directrix):

1x + (-5) y = 10 [likely given as x - 5y = 10] Standard Form: the standard form of a line is in the form Ax + By = C where A is a positive integer, and B, and C are integers.